# The Poisson boundary of discrete subgroups of semisimple Lie groups without moment conditions

### v1

@@ Abstract.
We show that the Poisson boundary of random walks of finite entropy on Zariski-dense discrete subgroups of semisimple Lie groups equals the Furstenberg boundary of the corresponding symmetric spaces equipped with the hitting measure, without assuming any moment condition on the random walk.
@@ 1. Introduction
Poisson boundaries capture algebraic and analytic properties of a group, for instance a countable group is amenable if and only if it admits a measure with trivial Poisson boundary [KV83, Ros81]. In many cases the group $G$ acts by isometries on a metric space $(X,d)$, the space $X$ is equipped with a natural topological boundary $\partial X$, and one of the central questions in the field has been the “identification problem”, that is whether the Poisson boundary coincides with $\partial X$, in the sense that the isomorphism above is realized by setting on $B=\partial X$ and $\nu$ to be the hitting measure (see e.g. [Kai96, Zhe22]).
@@
If $G$ is a semisimple Lie group and $\mu$ is absolutely continuous with respect to the Haar measure on $G$, then Furstenberg [Fur63b] showed that the Poisson boundary of $(G,\mu)$ can be identified with the space $B=G/P$, where $P$ is a minimal parabolic subgroup, which is known as the Furstenberg boundary. See [Bab06, Fur02] for a survey.
On the other hand, if the support of $\mu$ is countable, these results do not directly apply. Let now $\mu$ be a countably supported measure on a semisimple Lie group $G$, and suppose that the semigroup generated by the support of $\mu$ is a discrete subgroup $\Gamma$.
@@
If the measure $\mu$ has finite first moment, then Furstenberg [Fur63a] showed that almost every sample path converges in the boundary, and the drift exists and is positive. In this case, Ledrappier [Led85] showed the Poisson boundary is the Furstenberg boundary.
@@
Later, Kaimanovich [Kai00] proved that the Furstenberg boundary is the Poisson boundary under the weaker condition that $\mu$ has finite entropy and finite logarithmic moment. In order to do so, he devised two general criteria (the ray approximation and the strip approximation) to prove the identification of the Poisson boundary with the geometric boundary. This criteria have been applied in a wealth of contexts, especially to groups acting on spaces whose geometry is, in various ways, nonpositively curved. However, these criteria require a finite moment condition on the measure $\mu$ (finite first moment for ray approximation, finite logarithmic moment for the strip approximation).
@@
In this paper, we drop every moment condition, and identify the Poisson boundary for random walks on discrete subgroups of semisimple Lie groups assuming just that entropy is finite. Note that if $\mu$ has finite first moment, then the random walk converges almost surely in the visual compactification of the symmetric space $S=G/K$, where $K$ is a maximal compact subgroup; on the other hand, if one drops the moment condition, the random walk still converges, but in a weaker sense, i.e. in the Furstenberg-Satake compactification, whose boundary is $G/P$ [GR85]. Yet, this is sufficient to define a boundary map, hence it makes sense to ask whether the resulting topological boundary is the Poisson boundary.
Further, let us remark that finite entropy is necessary: in fact, one can find measures $\mu$, even on the free group, such that the Poisson boundary is larger than the geometric boundary [CF25]; one can adapt this counterexample to the context of discrete subgroups of Lie groups by e.g. embedding a free group as a discrete subgroup of $\textup{SL}(2,\mathbb{R})$ (in fact, such counterexamples exist on any finitely generated group of superpolynomial growth [CF25, Theorem 1.4]).
@@
Let $G$ be a semisimple, connected, Lie group with finite center and no compact factors, and let $\Gamma<G$ be a discrete, Zariski-dense subgroup. Let $\mu$ be a probability measure on $\Gamma$, with finite entropy, such that its support generates $\Gamma$ as a semigroup. Then the Furstenberg boundary $(G/P,\nu)$ with the hitting measure $\nu$ is the Poisson boundary for $(\Gamma,\mu)$.
@@
This result is the higher rank analog of the main theorem of [CFFT25], that for hyperbolic groups, and more generally acylindrically hyperbolic groups, establish the identification of the Poisson boundary with the Gromov boundary under finite entropy and without moment conditions.
In the footsteps of [CFFT25], we replace the strip approximation from [Kai00] by the so called pin-down approximation: namely, one identifies a sequence $(p_{n})$ of partitions of the path space, which can be interpreted as revealing some additional information on the walk, such that:
@@
    if $(B,\nu)$ is a $\mu$-boundary, revealing both the boundary point $\xi\in B$ and the outcome of $p_{n}$ determines, or “pins down”, the location $w_{n}$, with a subexponential error: $H_{B}(\alpha_{n}|p_{n})=o(n)$.
@@
    the total information contained in revealing $p_{n}$ is still small: $H(p_{n})=o(n)$.
@@
Then, the challenge becomes to construct carefully the partitions $(p_{n})$, by revealing certain information on the random walk (e.g., the distance from the origin). At this level, the strategy is similar to [CFFT25] (also used in [FS23]), but in this paper we need to introduce some new techniques.
First of all, in higher rank one needs to replace the limiting geodesic ray by a flat, and one cannot, differently from the rank one case, associate a flat to just one boundary point; thus, we need to run the random walk both forward and backward, and then consider the flat joining the boundary points of the two walks. Now, given that flats also grow polynomially, it is possible to estimate the conditional entropy by projecting onto the flats.
However, the most important difference is that in our previous work [CFFT25] we construct the partitions $(p_{n})$ by using the pivoting theory of Gouëzel; in particular, we keep track of certain times, called pivotal times, when the random walk is “aligned” with the limiting geodesic, and we cite the strong quantitative estimates on pivotal times provided by [Gou22]. Even though an analogy to pivoting theory in higher rank Lie groups has been established [Pén24], in this paper we do not attempt to use it.
Instead, we go a different route, adopting a softer approach, that would also provide a simpler proof in the rank 1 case. In this paper, we replace pivotal times by just times where the random walk is within some bounded distance of the flat; abundance of such times, that we call critical times, is guaranteed by a simple recurrence argument using the ergodicity of the shift, without the need for the strong exponential bounds given by pivoting theory (see Lemma 5.2).
@@ Acknowledgements
B.F. is partially supported by NSF grant DMS-2246727, J.F. is partially supported by NSF grant DMS-2348981 and G.T. is partially supported by NSERC grant RGPIN-2024-04324. We thank Vadim Kaimanovich for useful conversations.

### v2

@@ Abstract.
We consider random walks of finite entropy on Zariski-dense discrete subgroups of semisimple Lie groups and show that their Poisson boundary is identified with the Furstenberg boundary of the corresponding symmetric spaces equipped with the hitting measure. We do not assume any moment condition on the random walk and, in contrast with the previous rank $1$ proof, we do not rely on pivoting theory.
@@ 1. Introduction
Poisson boundaries capture algebraic and analytic properties of a group, for instance a countable group is amenable if and only if it admits a measure with trivial Poisson boundary [KV83, Ros81]. In many cases the group $G$ acts by isometries on a metric space $(X,d)$, the space $X$ is equipped with a natural topological boundary $\partial X$, and one of the central questions in the field has been the “identification problem”, that is whether the Poisson boundary coincides with $\partial X$, in the sense that the isomorphism above is realized by setting $B=\partial X$ and $\nu$ to be the hitting measure (see e.g. [Kai96, Sections 2.4, 2.8], [Zhe22, Section 1]).
@@
If $G$ is a semisimple Lie group and $\mu$ is absolutely continuous with respect to the Haar measure on $G$, then Furstenberg [Fur63b] showed that the Poisson boundary of $(G,\mu)$ can be identified with the space $B=G/P$, where $P$ is a minimal parabolic subgroup, which is known as the Furstenberg boundary. See [Fur02, Sections 2.4-2.6] and [Bab06, Sections 2.6, 6.2] for a survey.
On the other hand, if the support of $\mu$ is countable, these results and techniques do not directly apply; as observed in [Kai96, page 149], this is because the automorphism group of the corresponding Markov operator is not large enough to act transitively on the candidate Poisson boundary.
@@
Let now $\mu$ be a countably supported measure on a semisimple Lie group $G$, and suppose that the semigroup generated by the support of $\mu$ is a discrete subgroup $\Gamma$. If the measure $\mu$ has finite first moment, then Furstenberg [Fur63a] showed that almost every sample path converges in the boundary, and the drift exists and is positive. In this case, Ledrappier [Led85] and Kaimanovich [Kai85] showed that the Poisson boundary of the discrete group, i.e. of $(\Gamma,\mu)$, is identified with the Furstenberg boundary $G/P$ of the ambient Lie group $G$. This identification allows one to compare Lie groups with the lattices contained in them, yielding e.g. rigidity statements, such as that certain abstract groups cannot be lattices in certain Lie groups ([Fur67], [BF14]).
@@
Later, Kaimanovich [Kai00] extended this identification result, between the Poisson boundary of the discrete group and the Furstenberg boundary of the ambient Lie group, to the more general case that $\mu$ has finite entropy and finite logarithmic moment. In order to do so, he devised two general criteria to prove the identification of the Poisson boundary with the geometric boundary, known as the ray approximation and the strip approximation. These criteria have been applied in a wealth of contexts, especially to groups acting on spaces whose geometry is, in various ways, nonpositively curved: a few examples are [BL94, Kai94, CKW94, KM96, KM99, KW02, GM12, Hor16, MT18], among others. However, these criteria require some moment condition on the measure $\mu$ (finite first moment for ray approximation, finite logarithmic moment for the strip approximation) to control the geometry of the random walk.
@@
In this paper, we drop every moment condition, and identify the Poisson boundary for random walks on discrete subgroups of semisimple Lie groups assuming just that the entropy is finite. Note that if $\mu$ has finite first moment, then the random walk almost surely converges in the visual compactification of the symmetric space $S=G/K$, where $K$ is a maximal compact subgroup. On the other hand, if one drops the moment condition, the random walk still converges, but in a weaker sense, i.e. in the Furstenberg-Satake compactification, whose boundary is $G/P$ [GR85]. Yet, this is sufficient to define a boundary map, hence it makes sense to ask whether the resulting topological boundary is the Poisson boundary.
@@
Let $G$ be a semisimple, connected, Lie group with finite center and no compact factors, and let $\Gamma<G$ be a discrete, Zariski-dense subgroup. Let $\mu$ be a probability measure on $\Gamma$ with finite entropy, such that its support generates $\Gamma$ as a semigroup. Then the Furstenberg boundary $(G/P,\nu)$ with the hitting measure $\nu$ is the Poisson boundary for $(\Gamma,\mu)$.
@@
We note that finite entropy is a necessary assumption for our result and not just a technical artifact. Poisson boundaries exhibit radically different phenomena in the presence of finite entropy as opposed to the general case; for example, in infinite entropy there are measures $\mu$ with trivial boundary such that $\mu^{-1}$ has a nontrivial boundary ([Kai83], [Alp21]), examples of measures on products of groups $G\times H$ whose boundary is nontrivial despite the quotient boundaries on $G$ and $H$ being trivial ([Kai24], [Alp24]) and, most relevantly for the current project, it is in general impossible to identify a universal space as the Poisson boundary for any group of superpolynomial growth [CF25, Theorem 1.4]. In particular our result is false in general if we do not include the finite entropy assumption, as there are measures $\mu$, for instance on discrete subgroups of $\textup{SL}(2,\mathbb{R})$, for which the Poisson boundary is larger than the geometric boundary [CF25].
Theorem 1.1 is the higher rank analog of the main theorem of [CFFT25], that for hyperbolic groups, and more generally acylindrically hyperbolic groups, establish the identification of the Poisson boundary with the Gromov boundary under finite entropy and without moment conditions.
In the footsteps of [CFFT25](and also [FS23]), we replace the strip approximation from [Kai00] by what we call the pin-down approximation: namely, for any $\epsilon>0$ one identifies a sequence $(p_{n})$ of partitions of the path space, which can be interpreted as revealing some additional information on the walk, such that:
@@
    if $(B,\nu)$ is a $\mu$-boundary, revealing both the boundary point $\xi\in B$ and the outcome of $p_{n}$ determines, or “pins down”, the location $w_{n}$, with a subexponential error: this shows that the conditional entropy satisfies $H_{B}(w_{n}|p_{n})=o(n)$;
@@
    the total information contained in revealing $p_{n}$ is still small: $H(p_{n})\leq\epsilon n$.
@@
Then, the challenge becomes to construct carefully the partitions $(p_{n})$, by revealing certain information on the random walk (e.g., the distance from the origin), and at this step new techniques are required.
In our previous work [CFFT25], we made crucial use of the pivoting theory of Gouëzel [Gou22]. In fact, in order to construct the partitions $(p_{n})$ we kept track of certain times, called pivotal times, when the random walk is aligned with the limiting geodesic in a fairly strong sense. The existence and abundance of such times is there guaranteed by [Gou22], that shows strong exponential bounds on the probability of finding such pivots: such quantitative estimates were heavily used in [CFFT25]. Let us note that an analogue to pivoting theory in higher rank Lie groups has been established in the recent exciting work [Pén24], but it is not clear if it can be applied, and we do not use it in this paper. In fact, our main theorem (Theorem 1.1) is stated there as a conjecture ([Pén24, Conjecture 1.12]).
In the current paper, instead, we take a different route, avoiding pivoting theory and the corresponding exponential bounds altogether. We replace pivotal times by all times where the random walk is within some bounded distance of the flat; abundance of such times, that we call critical times, is guaranteed by a simple recurrence argument using the ergodicity of the shift, without the need for the strong exponential bounds given by pivoting theory (see Lemma 5.2). Let us note that this softer approach would also provide a simpler proof in the rank 1 case.
Moreover, the second difficulty is that in higher rank one needs to replace the limiting geodesic ray by a flat, and one cannot, in contrast with the rank one case, associate a flat to just one boundary point. Therefore, we need to run the random walk both forwards and backwards, and then consider the flat joining the boundary points of the two walks. The pin down partition is then constructed by replacing the distance along a geodesic ray by the projection of the random walk onto the flats, using the Cartan projection. The fact that flats are translates of abelian groups and have polynomial growth then makes it possible to estimate the conditional entropy, concluding the argument.
@@ Acknowledgements
B.F. is partially supported by NSF grant DMS-2246727, J.F. is partially supported by NSF grant DMS-2348981 and G.T. is partially supported by NSERC grant RGPIN-2024-04324. We thank Vadim Kaimanovich and Keivan Mallahi-Karai for useful conversations.